# Equal contribution    Project lead    Corresponding author

### v2

@@ 1 Introduction
Increasing model size and training tokens typically leads to performance improvements, as known as scaling laws, which have been extensively explored in the language and vision domains (Kaplan et al., 2020; Alabdulmohsin et al., 2022). However, such properties have not been thoroughly investigated in the time series domain. Assuming that scaling forecasting models with high-quality training data follows similar principles, several challenges remain: Dense versus sparse training. Most time series forecasting models compose of dense layers, which means each input time series tokens requires computations with all model parameters. While effective, this is computationally intensive. In contrast, sparse training with mixture-of-experts (MoE) is more flop-efficient per parameter and allows for scaling up model size with a fixed inference budget while giving better performance, as showcased on the right of Figure 1. However, optimizing a sparse, large-scale time series model faces another challenge of stability and convergency. Time series are highly heterogeneous (Woo et al., 2024; Dong et al., 2024), and selecting the appropriate model design and routing algorithm often involves a trade-off between performance and computational efficiency. Sparse solutions for time series foundation models have yet to be explored, leaving a significant gap in addressing these two challenges. While time series pre-training datasets are no longer a major bottleneck, most existing works (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024) have not extensively discussed their in-model data processing pipelines or mixing strategies. Answering this is particularly important, given that existing data archives are often noisy and largely imbalanced across domains.
On the other hand, most time series FMs face limitations in flexibility and generalizability. General-purpose forecasting is a fundamental capability, requiring a model to handle any forecasting problems, regardless of context lengths, forecasting horizons, input variables, and other properties such as frequencies and distributions. Meanwhile, achieving strong generalizability pushes the boundaries further that existing works often fail to meet simultaneously. For instance, Timer (Liu et al., 2024b) has limited native support for arbitrary output lengths, which may lead to truncated outputs, while Moment (Goswami et al., 2024) operates with a fixed input context length. Although Moirai (Woo et al., 2024) achieves universal forecasting, it depends on hardcoded heuristics in both the input and output layers.
@@ 2 Related Work
Time Series Forecasting. Deep learning models have become powerful tools for time series forecasting over the past decade, which can be broadly categorized into two types: (1) univariate models, such as DeepState (Rangapuram et al., 2018), DeepAR (Salinas et al., 2020), and N-BEATS (Oreshkin et al., 2020), which focus on modeling individual time series, and (2) multivariate models, which include both transformer-based approaches (Wen et al., 2023; Zhou et al., 2021; Nie et al., 2023; Liu et al., 2024a; Wang et al., 2024c; Chen et al., 2024) and non-transformer models (Sen et al., 2019; Jin et al., 2022; Wang et al., 2024b; Hu et al., 2024; Qi et al., 2024), designed to handle multiple time series simultaneously. While these models achieve competitive in-domain performance, many are task-specific and fall short in generalizability when applied to cross-domain data in few-shot or zero-shot scenarios.
@@
Sparse Deep Learning for Time Series. Deep learning models are often dense and over-parameterized (Hoefler et al., 2021), leading to increased memory and computational demands during both training and inference. However, sparse networks, such as mixture-of-experts models (Jacobs et al., 1991), which dynamically route inputs to specialized expert networks, have shown comparable or even superior generalization to dense models while being more efficient (Fedus et al., 2022; Riquelme et al., 2021). In time series research, model sparsification has received less attention, as time series models have traditionally been small in scale, with simple models like DLinear (Zeng et al., 2023) and SparseTSF (Lin et al., 2024) excelling in specific tasks prior to the advent of large-scale, general pre-training. The most relevant works on this topic include Pathformer (Chen et al., 2024), MoLE (Ni et al., 2024), and IME (Ismail et al., 2023). However, none of them delve into the scalability of foundation models with sparse structures. Besides, MoLE and IME are not sparse models, as input data is passed to all heads and then combined to make predictions.
@@ 3 Methodology
Figure 2: The architecture of Time-MoE, which is a decoder-only model. Given an input time series of arbitrary length, 1 we first tokenize it into a sequence of data points, 2 which are then encoded. These tokens are processed through $N$-stacked backbone layers, primarily consisting of causal multi-head self-attention and 3 sparse temporal mixture-of-expert layers. During training, 4 we optimize forecasting heads at multiple resolutions. For model inference, Time-MoE provides forecasts of flexible length by 5 dynamically scheduling these heads.
@@ 3.1 Time-MoE Overview
$s_{i,t}$ $=\operatorname{Softmax}_{i}\left(\mathbf{W}_{i}^{l}\mathbf{u}_{t}^{l}\right),$ (8)
where $\mathbf{W}_{i}^{l}\in\mathbb{R}^{1\times D}$ denotes the trainable parameters, and $N$ and $K$ respectively denote the numbers of non-shared experts and activated non-shared experts per MoE layer.
@@ 3.2.1 Time-300B Dataset
Training time series foundation models require extensive, high-quality data. However, well-processed large-scale datasets are still relatively scarce. Recent advancements have facilitated the collection of numerous time series datasets from various sources (Godahewa et al., 2021; Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024b). Nonetheless, data quality still remains a challenge, with prevalent issues such as missing values and invalid observations (Wang et al., 2024a) that can impair model performance and destabilize training. To mitigate these issues, we developed a streamlined data-cleaning pipeline (Appendix C) to filter and refine raw data, and constructed the largest open-access, high-quality time series data collection named Time-300B for foundation model pre-training. Time-300B is composed of a diverse range of publicly available datasets, spanning multiple domains such as energy, retail, healthcare, weather, finance, transportation, and web, as well as a portion of synthetic data to enhance the data quantity and diversity. Time-300B covers a wide spectrum of sampling frequencies from seconds to yearly intervals, and has over 300 billion time points after being processed by our data-cleaning pipeline, as summarized in Table 1.
@@
% 5.17 % 0.0001% 0.0001% 90.50 % 0.008 % 2.98% 0.69 % 0.58 % 0.006 % 100%
Table 2: A high-level summary of Time-MoE model configurations.
Time-MoEbase 12 12 8 2 384 1536 192 50 $\mathrm{M}$ 113 $\mathrm{M}$
Time-MoElarge 12 12 8 2 768 3072 384 200 $\mathrm{M}$ 453 $\mathrm{M}$
Time-MoEultra 36 16 8 2 1024 4096 512 1.1 $\mathrm{B}$ 2.4 $\mathrm{B}$
@@ 4 Main Results
Time-MoE consistently outperforms state-of-the-art forecasting models by large margins across six well-established benchmarks and settings (Appendix B). To ensure a fair comparison, we adhered to the experimental configurations from (Woo et al., 2024) for out-of-distribution forecasting and (Wu et al., 2023a) for in-distribution forecasting with a unified evaluation pipeline we developed. Specifically, we evaluate Time-MoE against 16 different baselines, representing state-of-the-art models in long-term forecasting. They are categorized into two groups: 1. zero-shot forecasting evaluation group, includes pre-trained foundation models such as Moirai (2024), TimesFM (2024), Moment (2024), and Chronos (2024); 2. in-distribution (full-shot) forecasting evaluation group, consists of modern time series models such as iTransformer (2024a), TimeMixer (2024b), TimesNet (2023a), PatchTST (2023), Crossformer (2023), TiDE (2023), DLinear (2023), and FEDformer (2022b).
@@ 4.1 Zero-shot Forecasting
Table 3: Full results of zero-shot forecasting experiments. A lower MSE or MAE indicates a better prediction. TimesFM, due to its use of Weather datasets in pretraining, is not evaluated on these two datasets and is denoted by a dash (--). Red: the best, Blue: the 2nd best.
@@
Setup. Time series foundation models have recently demonstrated impressive zero-shot learning capabilities (Liang et al., 2024). In this section, we conducted experiments on the six well-known long-term forecasting benchmarks for which datasets were not included in the pre-training corpora. We use four different prediction horizons, which are $\{96,192,336,720\}$, with the corresponding input time series lengths $\{512,1024,2048,3072\}$. The evaluation metrics adopt mean square error (MSE) and mean absolute error (MAE).
Results. Detailed results of zero-shot forecasting are in Table 3. Time-MoE achieves consistent state-of-the-art performances, improving a large margin as MSE reduction in average exceeding 20% over the other most competitive baselines. Importantly, as the model size scales (e.g., base $\rightarrow$ ultra), it continuously exhibits enhanced performance across all datasets, affirming the efficacy of scaling laws within our time series foundation models. Furthermore, in comparisons with robust baselines that have a similar number of activated parameters, Time-MoE demonstrates significantly superior performance. The largest models among the state-of-the-art baselines are Chronoslarge, Moment and Moirailarge. Compared to those models, Time-MoE achieved average MSE reductions of 23%, 30% and 11% respectively.
@@ 4.2 In-distribution Forecasting
Table 4: Full results of in-domain forecasting experiments. A lower MSE or MAE indicates a better prediction. Full-shot results besides Global Temp are obtained from (Liu et al., 2024a). Red: the best, Blue: the 2nd best.
@@ 4.3 Ablation Study
Table 5: Ablation studies. (Left) Average MSE for horizon-96 forecasting across six benchmarks, evaluated with different model components. (Right) Analysis of various multi-resolution forecasting configurations. Further details in Appendix D.1.
@@
Time-MoEbase 0.262 0.095 s/iter
@@ 4.4 Scalability Analysis
We save model checkpoints at intervals of every 20 billion time points during training, allowing us to plot performance traces for models of different sizes trained on various data scales. The right side of Figure 3 shows that models trained on larger datasets consistently outperform those trained on smaller datasets, regardless of model size. Our empirical results confirm that as both data volume and model parameters scale, sparse models demonstrate continuous and substantial improvements in performance, as well as achieve better forecasting accuracy compared to the dense counterparts under the same scales.
@@
Table 6: Comparison of BF16 and FP32 in terms of training and inference efficiency. Further details are provided in Table 13 of Appendix D.2. FA denotes flash-attention.
@@ Appendix A Further Related Work
In this section, we delve deeper into the related work on large time series models. Current research efforts in universal forecasting with time series foundation models can be broadly classified into three categories, as summarized in Table 8: (1) encoder-only models, such as Moirai (Woo et al., 2024) and Moment (Goswami et al., 2024), which employ masked reconstruction and have been pre-trained on datasets containing 27B and 1B time points, respectively, with model sizes reaching up to 385M parameters; (2) encoder-decoder models, exemplified by Chronos (Ansari et al., 2024), which offers pre-trained models at four scales, with up to 710M parameters; and (3) decoder-only models, including TimesFM (Das et al., 2024), Lag-Llama (Rasul et al., 2023), and Timer (Liu et al., 2024b), with the largest models containing up to 200M parameters. In contrast to these dense models, Time-MoE introduces a scalable, unified architecture with a sparse mixture-of-experts design, optimized for larger time series forecasting models while reducing inference costs. Trained on our Time-300B dataset, comprising over 300B time points, Time-MoE is scaled to 2.4B parameters for the first time. It outperforms existing models with the same number of activated parameters, significantly enhancing both model efficiency and forecasting precision, while avoiding limitations such as fixed context lengths or hardcoded heuristics.
@@
* Depend on the way of calculation according to the original paper.
@@ Appendix C Processed Data Archive
Going beyond previous work (Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024b), we organized a comprehensive large-scale time series dataset from a vast collection of complex raw data. To ensure data quality, we addressed issues by either imputing missing values or discarding malformed time series. Inspired by data processing techniques from large language models (Penedo et al., 2023; Computer, 2023; Jin et al., 2024), we developed a fine-grained data-cleaning pipeline specifically designed for time series data:
@@
Algorithm 2 Sample code of Data-cleaning Pipline

### v3

@@ 1 Introduction
Increasing model size and training tokens typically leads to performance improvements, as known as scaling laws, which have been extensively explored in the language and vision domains (Kaplan et al., 2020; Alabdulmohsin et al., 2022). However, such properties have not been thoroughly investigated in the time series domain (Yao et al., 2025). Assuming that scaling forecasting models with high-quality training data follows similar principles, several challenges remain: Dense versus sparse training. Most time series forecasting models compose of dense layers, which means each input time series tokens requires computations with all model parameters. While effective, this is computationally intensive. In contrast, sparse training with mixture-of-experts (MoE) is more flop-efficient per parameter and allows for scaling up model size with a fixed inference budget while giving better performance, as showcased on the right of Figure 1. However, optimizing a sparse, large-scale time series model faces another challenge of stability and convergency. Time series are highly heterogeneous (Woo et al., 2024; Dong et al., 2024), and selecting the appropriate model design and routing algorithm often involves a trade-off between performance and computational efficiency. Sparse solutions for time series foundation models have yet to be explored, leaving a significant gap in addressing these two challenges. While time series pre-training datasets are no longer a major bottleneck, most existing works (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024) have not extensively discussed their in-model data processing pipelines or mixing strategies. Answering this is particularly important, given that existing data archives are often noisy and largely imbalanced across domains.
On the other hand, most time series FMs face limitations in flexibility and generalizability. General-purpose forecasting is a fundamental capability, requiring a model to handle any forecasting problems, regardless of context lengths, forecasting horizons, input variables, and other properties such as frequencies and distributions. Meanwhile, achieving strong generalizability pushes the boundaries further that existing works often fail to meet simultaneously. For instance, Timer (Liu et al., 2024d) has limited native support for arbitrary output lengths, which may lead to truncated outputs, while Moment (Goswami et al., 2024) operates with a fixed input context length. Although Moirai (Woo et al., 2024) achieves universal forecasting, it depends on hardcoded heuristics in both the input and output layers.
@@ 2 Related Work
Time Series Forecasting. Deep learning models have become powerful tools for time series forecasting over the past decade, which can be broadly categorized into two types: (1) univariate models, such as DeepState (Rangapuram et al., 2018), DeepAR (Salinas et al., 2020), and N-BEATS (Oreshkin et al., 2020), which focus on modeling individual time series, and (2) multivariate models, which include both transformer-based approaches (Wen et al., 2023; Zhou et al., 2021; Nie et al., 2023; Liu et al., 2024b; Wang et al., 2024c; Chen et al., 2024) and non-transformer models (Sen et al., 2019; Jin et al., 2022; Wang et al., 2024b; Hu et al., 2024; Qi et al., 2024), designed to handle multiple time series simultaneously. While these models achieve competitive in-domain performance (Wang et al., 2025), many are task-specific and fall short in generalizability when applied to cross-domain data in few-shot or zero-shot scenarios.
@@
Sparse Deep Learning for Time Series. Deep learning models are often dense and over-parameterized (Hoefler et al., 2021), leading to increased memory and computational demands during both training and inference. However, sparse networks, such as mixture-of-experts models (Jacobs et al., 1991), which dynamically route inputs to specialized expert networks, have shown comparable or even superior generalization to dense models while being more efficient (Fedus et al., 2022; Riquelme et al., 2021). In time series research, model sparsification has received relatively less attention, as time series models have traditionally been small in scale, with simple models like DLinear (Zeng et al., 2023) and SparseTSF (Lin et al., 2024) excelling in specific tasks prior to the advent of large-scale, general pre-training. The most relevant works on this topic include Pathformer (Chen et al., 2024), MoLE (Ni et al., 2024), and IME (Ismail et al., 2023). However, none of them delve into the scalability of foundation models with sparse structures. Besides, MoLE and IME are not sparse models, as input data is passed to all heads and then combined to make predictions.
@@ 3 Methodology
Figure 2: The architecture of Time-MoE, which is a decoder-only model. Given an input time series of arbitrary length, 1 we first tokenize it into a sequence of data points, 2 which are then encoded. These tokens are processed through $N$-stacked backbone layers, primarily consisting of causal multi-head self-attention and 3 sparse temporal mixture-of-expert layers. During training, 4 we optimize forecasting heads at multiple resolutions. For model inference, Time-MoE provides forecasts of flexible length by 5 dynamically scheduling these heads. Details about the causal multi-head self-attention are in Appendix B and illustrated in Figure 5.
@@ 3.1 Time-MoE Overview
$s_{i,t}$ $=\operatorname{Softmax}_{i}\left(\mathbf{W}_{i}^{l}\mathbf{\bar{u}}_{t}^{l}\right),$ (8)
where $\mathbf{W}_{i}^{l}\in\mathbb{R}^{1\times D}$ denotes the trainable parameters, and $N$ and $K$ respectively denote the numbers of non-shared experts and activated non-shared experts per mixture-of-experts layer.
@@ 3.2.1 Time-300B Dataset
Training time series foundation models require extensive, high-quality data. However, well-processed large-scale datasets are still relatively scarce. Recent advancements have facilitated the collection of numerous time series datasets from various sources (Godahewa et al., 2021; Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024d; a). Nonetheless, data quality still remains a challenge, with prevalent issues such as missing values and invalid observations (Wang et al., 2024a) that can impair model performance and destabilize training. To mitigate these issues, we developed a streamlined data-cleaning pipeline (Appendix C) to filter and refine raw data, and constructed the largest open-access, high-quality time series data collection named Time-300B for foundation model pre-training. Time-300B is composed of a diverse range of publicly available datasets, spanning multiple domains such as energy, retail, healthcare, weather, finance, transportation, and web, as well as a portion of synthetic data to enhance the data quantity and diversity. Time-300B covers a wide spectrum of sampling frequencies from seconds to yearly intervals, and has over 300 billion time points after being processed by our data-cleaning pipeline, as summarized in Table 1.
@@
Percent5.17 % 0.0001% 0.0001% 90.50 % 0.008 % 2.98% 0.69 % 0.58 % 0.006 % 100%
@@ 3.2.3 Model Configurations and Training Details
Table 2: A high-level summary of Time-MoE model configurations.
Time-MoEbase 12 12 8 2 384 1536 192 50 $\mathrm{M}$ 113 $\mathrm{M}$
Time-MoElarge 12 12 8 2 768 3072 384 200 $\mathrm{M}$ 453 $\mathrm{M}$
Time-MoEultra 36 16 8 2 1024 4096 512 1.1 $\mathrm{B}$ 2.4 $\mathrm{B}$
@@ 4 Main Results
Time-MoE consistently outperforms state-of-the-art models by large margins across 6 well-established benchmarks and settings (Appendix B). To ensure a fair comparison, we adhered to the configurations from (Woo et al., 2024) for out-of-distribution forecasting and (Wu et al., 2023a) for in-distribution forecasting with a unified evaluation pipeline we developed. Specifically, we evaluate Time-MoE against 16 different baselines, representing state-of-the-art models in long-term forecasting. They are categorized into two groups: (1) zero-shot forecasting evaluation group, includes pre-trained foundation models such as Moirai (2024), TimesFM (2024), Moment (2024), and Chronos (2024); (2) in-distribution (full-shot) forecasting evaluation group, consists of up-to-date models such as iTransformer (2024b), TimeMixer (2024b), TimesNet (2023a), PatchTST (2023), Crossformer (2023), TiDE (2023), DLinear (2023),FEDformer (2022b). We also include addition comparisons with Timer  (2024d), TFT (2021), and N-BEATS (2020) in Appendix D.3.
@@ 4.1 Zero-shot Forecasting
Table 3: Full results of zero-shot forecasting experiments. A lower MSE or MAE indicates a better prediction. TimesFM, due to its use of Weather datasets in pretraining, is not evaluated on this dataset and is denoted by a dash (--). Red: the best, Blue: the 2nd best.
@@
Setup. Time series foundation models have recently demonstrated impressive zero-shot learning capabilities (Liang et al., 2024; Liu et al., 2024c). In this section, we conducted experiments on the six well-known long-term forecasting benchmarks for which datasets were not included in the pre-training corpora. We use four different prediction horizons, which are $\{96,192,336,720\}$, with the corresponding input time series lengths $\{512,1024,2048,3072\}$. The evaluation metrics adopt mean square error (MSE) and mean absolute error (MAE).
Results. Detailed results of zero-shot forecasting are in Table 3. Time-MoE achieves consistent state-of-the-art performances, improving a large margin as MSE reduction in average exceeding 20% over the other most competitive baselines. Importantly, as the model size scales (e.g., Time-MoEbase $\rightarrow$ Time-MoEultra), it continuously exhibits enhanced performance across all datasets, affirming the efficacy of scaling laws within our time series foundation models. Furthermore, in comparisons with robust baselines that have a similar number of activated parameters, Time-MoE demonstrates significantly superior performance. The largest models among the state-of-the-art baselines are Chronoslarge, Moment and Moirailarge. Compared to those models, Time-MoE achieves average MSE reductions of 23%, 30% and 11% respectively.
@@ 4.2 In-distribution Forecasting
Table 4: Full results of in-domain forecasting experiments. A lower MSE or MAE indicates a better prediction. Full-shot results besides Global Temp are obtained from (Liu et al., 2024b). Red: the best, Blue: the 2nd best.
@@ 4.3 Ablation Study
Table 5: Ablation studies. (Left) Average MSE for horizon-96 forecasting across six benchmarks, evaluated with different model components. (Right) Analysis of various multi-resolution forecasting configurations. More details are in Appendix D.1.
@@
Time-MoEbase w/ {1,8,32,64} 0.262 0.095 s/iter
@@ 4.4 Scalability Analysis
We save model checkpoints at intervals of every 20 billion time points during training, allowing to plot performance traces for models of different sizes trained on various data scales. The right side of Figure 3 shows that models trained on larger datasets consistently outperform those trained on smaller datasets, regardless of model size. Our empirical results confirm that as both data volume and model parameters scale, sparse models demonstrate continuous and substantial improvements in performance, as well as achieve better forecasting accuracy compared to the dense counterparts under the same scales.
@@
Table 6: Comparison of BF16 and FP32 in terms of training and inference efficiency. FA denotes flash-attention. More details are in Table 13 of Appendix D.2.
@@ Appendix A Further Related Work
In this section, we delve deeper into the related work on large time series models. Current research efforts in universal forecasting with time series foundation models can be broadly classified into three categories, as summarized in Table 8: (1) encoder-only models, such as Moirai (Woo et al., 2024) and Moment (Goswami et al., 2024), which employ masked reconstruction and have been pre-trained on datasets containing 27B and 1B time points, respectively, with model sizes reaching up to 385M parameters; (2) encoder-decoder models, exemplified by Chronos (Ansari et al., 2024), which offers pre-trained models at four scales, with up to 710M parameters; and (3) decoder-only models, including TimesFM (Das et al., 2024), Lag-Llama (Rasul et al., 2023), and Timer (Liu et al., 2024d), with the largest models containing up to 200M parameters. The concurrent work, Moirai-MoE (Liu et al., 2024a), includes up to 935M parameters but with a different expert and routing design. In contrast to these models, Time-MoE introduces a scalable, unified architecture with a sparse mixture-of-experts design, optimized for larger time series forecasting models while reducing inference costs. Trained on our Time-300B dataset, comprising over 300B time points, Time-MoE is scaled to 2.4B parameters for the first time. It outperforms existing models with the same number of activated parameters, significantly enhancing both model efficiency and forecasting precision, while avoiding limitations such as fixed context lengths or hardcoded heuristics.
@@
* Depend on the way of calculation according to the original paper. indicates the total of the context and prediction lengths.
@@ Appendix B Implementation Details
Technical Details.
Our mixture-of-experts layer consists of one shared expert and several isolated experts, each represented by a feedforward network that is smaller than the standard FFN employed in dense models. In the formulation from Equations 5 to 8, $\text{FFN}_{N+1}$ denotes the shared expert, while $\text{FFN}_{1}$ to $\text{FFN}_{N}$ correspond to the isolated experts. The weight $g_{N+1,t}$ associated with the shared expert for token $t$ is normalized using the $\operatorname{Sigmoid}$ function. In contrast, the weight $g_{i,t}$ for the $i$-th isolated expert of token $t$ is normalized using the $\operatorname{Softmax}$ function. Furthermore, we retain only the top-$k$ largest scores among the isolated experts and set the remaining scores to zero.
Figure 5: Causal attention layer.
To prevent routing collapse among experts, we adopt the strategy proposed by  (Fedus et al., 2022), incorporating an auxiliary loss to ensure balanced expert load. The key aspect of this method is to penalize experts with high gating scores. This helps prevent a scenario where stronger experts, being exposed to more tokens, become even stronger while weaker experts continue to fall behind. The mathematical formulation is presented in Equation 10, where $f_{i}$ represents the fraction of tokens assigned to expert $i$, and $r_{i}$ denotes the proportion of router probability allocated to expert $i$. If one expert is assigned too many tokens and achieves a higher routing score, it will incur a correspondingly higher loss.
@@ Appendix C Processed Data Archive
Going beyond the previous work (Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024d), we organized a comprehensive large-scale time series dataset from a vast collection of complex raw data. We utilize the missing value ratio and the invalid observation ratio as metrics to assess the quality of the dataset. These two metrics can effectively identify data issues caused by the instability of data collection and artificially imputed values. The missing value ratio is defined as the proportion of ‘nan’ and ‘inf’ values present in the time series. Meanwhile, the invalid observation ratio refers to the maximum proportion of zeros in the first- or second-order differences of the time series. To address these issues and drawing inspiration from the data processing techniques of large language models (Penedo et al., 2023; Computer, 2023; Jin et al., 2024), we developed a fine-grained data-cleaning pipeline specifically designed for time series data:
@@
Algorithm 2 Code Snippet of Data-cleaning Pipline